# Sana: Efficient High-Resolution Image Synthesis with Linear Diffusion Transformers

### v2

@@ 3.1 Data Curation and Blending
Multi-Caption Auto-labelling Pipeline: For each image, whether or not it contains an original prompt, we will use four VLMs to label it: VILA-3B/13B (Lin et al., 2024), InternVL2-8B/26B (Chen et al., 2024c). Multiple VLMs can make the caption more accurate and more diverse.
@@ 4 On-device Deployment
Table 5 shows the speed comparison before and after our deployment optimization on a customer-grade edge device with a 4090 GPU. For generating a 1024px image, our optimized implementation achieves 2.4 speedup, taking only 0.37 seconds, while maintaining almost lossless image quality.
@@ Appendix A Full Related Work
Efficient Diffusion Transformers. The introduction of Diffusion Transformers (DiT) (Peebles & Xie, 2023) marked a significant shift in image generation models, replacing the traditional U-Net architecture with a transformer-based approach. This innovation paved the way for more efficient and scalable diffusion models. Building upon DiT, PixArt- (Chen et al., 2024b) extended the concept to text-to-image generation, demonstrating the versatility of transformer-based diffusion models. Stable Diffusion 3 (SD3) (Esser et al., 2024) further advanced the field by proposing the Multi-modal Diffusion Transformer (MM-DiT), which effectively integrates text and image modalities. More recently, Flux (Labs, 2024) showcased the potential of DiT architectures in high-resolution image generation by scaling up to 12B parameters. In addition, earlier works like CAN (Cai et al., 2024) explored linear attention mechanisms in class-condition image generation.
@@ Appendix C More Results
Finding: Zero-shot Language Transfer Ability. As shown in Figure 13, we were surprised to find that by using Gemma-2 as the text encoder and Chinese/Emoji expressions as text prompts; our Sana can also understand and generate corresponding images. Note that we filter out all prompts other than English during training, so the zero-shot generation capability of Chinese/Emoji is brought by Gemma-2.
@@
Detailed Speed Comparison of Text Encoder. In Table 13, we present a comparison of the latency and parameters for various T5 models alongside the Gemma models. Notably, the Gemma-2B model exhibits a similar latency to T5-large while significantly increasing the model size. This enhancement in model size is a key factor in achieving improved capabilities with greater efficiency.

### v3

@@ 3.1 Data Curation and Blending
Multi-Caption Auto-labelling Pipeline: For each image, whether or not it contains an original prompt, we will use four VLMs to label it: VILA-3B/13B (Lin et al., 2024), InternVL2-8B/26B (Chen et al., 2024d). Multiple VLMs can make the caption more accurate and more diverse.
@@ 4 On-device Deployment
Table 5 shows the speed comparison before and after our deployment optimization on a laptop GPU. For generating a 1024px image, our optimized implementation achieves 2.4 speedup, taking only 0.37 seconds, while maintaining almost lossless image quality.
@@ Appendix A Full Related Work
Efficient Diffusion Transformers. The introduction of Diffusion Transformers (DiT) (Peebles & Xie, 2023) marked a significant shift in image generation models, replacing the traditional U-Net architecture with a transformer-based approach. This innovation paved the way for more efficient and scalable diffusion models. Building upon DiT, PixArt- (Chen et al., 2024b) extended the concept to text-to-image generation, demonstrating the versatility of transformer-based diffusion models. Stable Diffusion 3 (SD3) (Esser et al., 2024) further advanced the field by proposing the Multi-modal Diffusion Transformer (MM-DiT), which effectively integrates text and image modalities. More recently, Flux (Labs, 2024) showcased the potential of DiT architectures in high-resolution image generation by scaling up to 12B parameters. In addition, earlier works like CAN (Cai et al., 2024) and DiG (Zhu et al., 2024) explored linear attention mechanisms in class-condition image generation. Several works are also related to modifying the model configuration, e.g., diffusion without attention (Yan et al., 2024; Teng et al., 2024) and cascade model structures (Pernias et al., 2023; Ren et al., 2024; Tian et al., 2024).
@@ Appendix B More Implementation Details
2K/4K fine-tuning. We reintroduce Positional Encoding (PE) to improve performance during the fine-tuning of 2K models on top of 1K models. For fine-tuning 4K models based on 2K models, we apply the same PE interpolation strategy used in Chen et al. (2024a). The training process with the addition of positional encoding (PE) converges remarkably quickly, typically within just 10K iterations.
@@ Appendix C More Results
Comparison Between Autoencoders Figure 12 illustrates the visual differences between the original images and the reconstructions generated by two distinct models: our DC-AE-F32C32 (Chen et al., 2024c) and SDXL’s VAE-F8C4. Both models deliver reconstructions nearly indistinguishable from the original images, showcasing their powerful encoding and decoding capabilities.
@@
Finding: Zero-shot Language Transfer Ability. As shown in Figure 14, we were surprised to find that by using Gemma-2 as the text encoder and Chinese/Emoji expressions as text prompts; our Sana can also understand and generate corresponding images. Note that we filter out all prompts other than English during training, so Gemma-2 brings the zero-shot generation capability of Chinese/Emoji.
@@
Detailed Speed Comparison of Text Encoder. In Table 13, we compare the latency and parameters for various T5 models alongside the Gemma models. Notably, the Gemma-2B model exhibits a similar latency to T5-large while significantly increasing the model size. This enhancement in model size is a key factor in achieving improved capabilities with greater efficiency.